OpenStudy (fuzzyball):

The value of y varies directly with the value of x, and y = –8 when x = 4. Which equations represent the relationship between y and x? Choose exactly two answers that are correct. A. x/y = 1/2 B. y/x = -2 C. -2x = y D. x=2y WILL MEDAL

OpenStudy (fuzzyball):

@MissSmartiez

OpenStudy (fuzzyball):

@welshfella

OpenStudy (fuzzyball):

@Loser66

OpenStudy (fuzzyball):

@Mehek14

OpenStudy (fuzzyball):

@Mehek14

OpenStudy (misssmartiez):

I am here :D!

OpenStudy (misssmartiez):

1: Turn all isolated variables into -8 and 4. 2: Solve: A. 4/-8 = 1/2 B. -8/4 = -2 C. -2 * 4 = -8 D. 4=2 * -8

OpenStudy (misssmartiez):

 D is false